OpenStudy (anonymous):

PLz help ASAp gold medal will be awarded. if g(x)=2(x-4), find the value of x if G(x)=20

OpenStudy (doggy):

g(x) = 2(x-4) 20 = 2x-8 20+8=2x 28=2x 28/2=x 14=x

OpenStudy (anonymous):

oh thank you so much i didnt know u do math to

OpenStudy (doggy):

no problem :3